# Open-Sora 2.0: Training a Commercial-Level Video Generation Model in \$200k

### v1

@@ 1 Introduction
The past year has witnessed an explosion of video generation models. Since the emergence of OpenAI’s Sora [undefc] in February 2024, a series of video generation models—either open-source [undefap, undefu, undefz, undefr, undefo] or proprietary [undefad, undefag, undefy]—have appeared at an unprecedented pace, each striving to achieve "Sora-level" generation quality. While the quality of generated videos continues to improve, there is a clear trend toward rapid growth in model size, data quantity, and computing resources. Following the success of scaling large language models (LLMs) [undefn, undefq], researchers are now applying similar scaling principles [undefr] to video generation, converging on similar model architectures and training techniques.
In this report, however, we show that a top-performing video generation model can be trained at a highly controlled cost, offering new insights into cost-effective training and efficiency optimization.. We build Open-Sora 2.0, a commercial-level video generation model trained for only \$200k. As shown in Table. 4, the training cost is 5-10 times lower than comparable models like MovieGen [undefad] and Step-Video-T2V [undefz] based on a fair comparison. This remarkable cost efficiency stems from our joint optimization of data curation, training strategy, and AI infrastructure—all of which are detailed in the following sections.
We show the human preference evaluation of Open-Sora 2.0 and other global leading video generation models in Figure 1. The models for comparison include proprietary models like Runway Gen-3 Alpha [undefag], Luma Ray2 [undefy] and open-source models like HunyuanVideo [undefr] and Step-Video-T2V [undefz]. We evaluate these models in three important aspects: 1) visual quality, 2) prompt adherence, and 3) motion quality. Despite the low cost of our model, it outperforms these top-performing models in at least two aspects out of the three, demonstrating its strong potential for commercial deployment.
@@ 2.1.1 Preprocessing
Then, we detect continuous shots within the raw video and segment it into short clips based on the detection result. For shot detection, we apply the libavfilter from FFmpeg [undefk] to calculate the scene scores, which quantify visual differences among frames. The shot boundaries are identified at points where the change in scene scores exceeds an empirically determined threshold.
@@ 2.1.2 Score Filtering
We assess the aesthetic quality of a sample using the CLIP+MLP aesthetic score predictor [undefah]. For a video sample, we extract the first, middle, and last frame, compute their individual scores, and take the average as the final aesthetic score.
@@
The motion intensity of a video is measured using the VMAF motion score from FFmpeg’s libavfilter library [undefk]. Videos with extremely low or excessively high motion scores are filtered out to maintain a balanced motion range.
@@
To evaluate image clarity, we apply the Laplacian operator from OpenCV [undefb]. A sample is considered blurry if the variance of the Laplacian image falls below a configurable threshold. For video samples, we extract five uniformly spaced frames and adopt a majority voting approach to determine blurriness.
@@
We detect text bounding boxes using PaddleOCR [undefa] and compute the total area of bounding boxes with a confidence score above 0.7. Images or videos containing excessive text are discarded to avoid unwanted overlays.
@@
We detect camera jitter using Shot Boundary Detection from the PySceneDetect library [undefe]. A video is identified as having camera jitter if the average frame-to-frame change exceeds a predefined threshold.
@@ 2.2 Data Annotation
For captioning, we utilize the open-source vision-language model LLaVA-Video [undefao] to annotate 256px videos. We prompt the model to focus on six aspects for a detailed and comprehensive caption, which are 1) main subjects; 2) subjects’ actions; 3) background and environment; 4) lighting condition and atmosphere; 5) camera movement; 6) video style, such as realistic, cinematic, 3D, animation, and so on. For high-resolution 768px training data, we leverage a stronger proprietary model Qwen 2.5 Max [undefak] to generate more accurate and semantically aligned captions. We find that Qwen 2.5 Max produces fewer hallucinations and delivers better semantic consistency than LLaVA-Video.
@@ 3 Model Architecture
Open Sora 1.2 [undefap] $4\times 8\times 8$ 192 4 $\checkmark$ 0.161 27.504 0.756
StepVideo VAE [undefz] $8\times 16\times 16$ 96 64 $\checkmark$ 0.082 28.719 0.818
HunyuanVideo VAE [undefr] $4\times 8\times 8$ 48 16 $\checkmark$ 0.046 30.240 0.856
@@
We begin by leveraging the open-source HunyuanVideo VAE [undefr] as the initial autoencoder for our model. To further reduce both training and inference costs, we develop a video autoencoder with deep compression (Video DC-AE, named after [undefh]) that improves efficiency while maintaining high reconstruction fidelity.
@@
[undefh] introduce DC-AE as an effective approach to achieve higher downsampling ratios while maintaining reconstruction quality. While the original DC-AE is primarily optimized for image encoding, we extend its architecture to support video encoding by incorporating temporal compression mechanisms.
As shown in Figure 5(a), our Video DC-AE encoder consists of three residual blocks followed by three EfficientViT blocks [undefd], with the decoder adopting a symmetrical structure. The first five encoder blocks and the last five decoder blocks function as downsampling and upsampling layers, respectively.
@@
Moreover, we follow DC-AE to introduce special residual blocks to connect the downsample and upsample blocks. This is because [undefh] identifies gradient propagation issues in these blocks, such that without the residuals, it is especially difficult to train high-compression autoencoders.
@@
We train Video DC-AE from scratch and evaluate its reconstruction quality, with results summarized in Table 1. Our model is compared against OpenSora 1.2 VAE, the recently open-sourced StepVideo VAE [undefz], and HunyuanVideo VAE [undefr]. The results demonstrate that Video DC-AE achieves competitive performance, with only minor degradation in LPIPS [undefan] scores compared to the best-performing HunyuanVideo VAE, while maintaining strong PSNR and SSIM scores.
@@ 3.2 DiT Architecture
To achieve higher video quality, we employ full attention to capture long-range dependencies effectively. After encoding the input through the autoencoder, we patchify the latent representations to enhance computational efficiency and improve model learning. Sana [undefal] suggests that a patch size of 1 yields better training stability and finer details in video generation. With Video DC-AE’s high compression ratio, we can afford to reduce the patch size to 1 (i.e., no patch at all), whereas a patch size of 2 is still required when the HunyuanVideo autoencoder is used in our generation model.
Inspired by FLUX’s MMDiT [undefs], we employ a hybrid transformer architecture that incorporates both dual-stream and single-stream processing blocks. In the dual-stream blocks, text and video information are processed separately to facilitate more effective feature extraction within each modality. Subsequently, single-stream blocks integrate these features to facilitate effective cross-modal interactions. To further enhance the model’s ability to capture spatial and temporal information, we apply 3D RoPE (Rotary Position Embedding) [undefaj], which extends traditional positional encoding to three-dimensional space, allowing the model to better represent motion dynamics across time.
For text encoding, we leverage T5-XXL [undefi] and CLIP-Large [undefae], two high-capacity pretrained models known for their strong semantic understanding. T5-XXL captures complex textual semantics, while CLIP-Large improves alignment between text and visual concepts, leading to more accurate prompt adherence in video generation.
@@ 4.1 Efficient Training Strategy
Prior research [undefaa, undefo, undefap] has demonstrated that pretraining on image datasets can significantly accelerate video model training. To avoid the high cost of training an image model from scratch, we leverage an open-source solution. Specifically, we adopt Flux [undefs], a state-of-the-art text-to-image model with 11 billion parameters, which provides sufficient capacity to generate high-quality videos. We initialize our text-to-video model using Flux and empirically find that, despite it being a distilled model, this initialization is effective for further training.
@@
Inspired by the efficient image training strategies from PixArt [undefg], we hypothesize that high-quality video data can substantially enhance training efficiency. Consequently, we curate a high-quality subset from a large-scale dataset for low-resolution training. For high-resolution fine-tuning, we impose stricter selection criteria to ensure superior video quality.
@@
With our proposed approach, we successfully constrain the one-time training cost to \$200K, as detailed in Table 3. We estimate the training costs of Movie Gen [undefad] and Step-Video-T2V [undefz] based on publicly available information, as summarized in Table 4. Our model achieves 5–10× lower training costs. We hope that these insights will contribute to reducing the training cost of high-quality video generation models in future research.
Model #GPUs GPU Hours Cost (Single Run)
Movie Gen [undefad] 6144 1.25M* \$2.5M*
Step-Video-T2V [undefz] 2992* 500k* \$1M*
Open Sora 2.0 224 100k \$200k
Table 4: Training Cost Comparison of Different Video Generation Models. Values marked with * are estimated based on publicly available information. Our Open-Sora 2.0 is trained at 5–10× lower training cost.
@@ 4.2 Training Setting
Our training setup is largely based on our previous iteration, Open-Sora 1.2 [undefap]. We adopt flow matching [undefv] as our primary training objective and utilize the AdamW [undefx] optimizer with $\beta$ values of (0.9, 0.999) and an $\epsilon$ value of $1\times 10^{-15}$. No weight decay is applied. The learning rate is set to $5\times 10^{-5}$ for the first 40k steps of Stages 1 and 2, followed by a decay to $3\times 10^{-5}$ for the final 45k steps. For Stage 3, the learning rate remains at $5\times 10^{-5}$, while in Stage 3, it is reduced to $1\times 10^{-5}$. Gradient norm clipping is applied with a threshold of 1.
@@
Our flow matching approach is similar to that used in Stable Diffusion 3 [undefj]. We denote the video latent as $X_{0}$, and Gaussian noise $X_{1}\sim\mathcal{N}(0,1)$. The model $f_{\theta}$ takes as input an interpolated latent $X_{t}=(1-t)X_{0}+tX_{1}$ and is trained to predict the velocity component $X_{0}-X_{1}$. The corresponding loss function is formulated as:
@@
Following the methodology in Open-Sora 1.2 [undefap], we adopt multi-bucket training to efficiently handle videos of varying frame counts, resolutions, and aspect ratios within the same batch. This strategy optimizes GPU utilization by dynamically assigning batch sizes based on video characteristics. The specific batch sizes used for different training configurations are detailed in Table 5 and Table 6.
@@ 4.3 High Compression Autoencoder Adaptation
For instance, in a 5-second, 24fps, 768px video, the HunyuanVideo VAE with a $4\times 8\times 8$ compression in combination of generation model’s patch size $1\times 2\times 2$ ($D_{\text{token}}=1024$) reduces the token count to 76K, whereas the Video DC-AE with $4\times 32\times 32$ compression and patch size of $1\times 1\times 1$ has a $D_{\text{token}}$ of 4096, effectively reduces the token count to 19K. As shown in Figure 7, this compression yields a $5.2\times$ speedup in training throughput. Additionally, inference speeds improve by over $10\times$, making it a highly efficient approach.
@@
With the growing interest in video generation models, we observe that despite variations in architecture, downsampling ratios, number of channels, and training losses, video autoencoders trained with similar information downsampling ratios exhibit comparable performance (evaluated at 256px resolution) [undefh]. We define the information downsampling ratio, $D_{\text{info}}$ as follows:
@@
However, training a deep video compression network is challenging. Prior work [undefh] identifies gradient propagation issues in downsample and upsample blocks. To address this, they introduce residual connections in these blocks using pixel shuffle and unshuffle operations. We verify that training the HunyuanVideo VAE architecture with residual connections improves performance. Building on this, we incorporate temporal downsampling into DC-AE to construct Video DC-AE.
@@
While training throughput for the diffusion model improves, we find that high-compression autoencoders—especially those with larger latent channel dimensions—slow down convergence. Prior studies [undefam, undefal] show that when training image diffusion models on AEs with better reconstruction but higher latent dimensionality, the generation quality often degrades. Additionally, Theorem A.7 in [undeff] suggests that, under strong assumptions, increasing the channel size by $k$ requires $k^{5}$ more data to achieve comparable performance. Meanwhile, [undefam] demonstrates that introducing a distillation loss between a pretrained image foundation model and the VAE latents can significantly accelerate diffusion training.
@@
    Latent Space Distillation: After training Video DC-AE, we apply a distillation loss to align the third-layer latents with DINOv2 [undefac], leveraging their similar latent shapes.
@@
    Tiling in video encoding: High compression AEs trained at low resolutions experience performance degradation when reconstructing high-resolution videos [undefh]. Although we could have fine-tuned the Video DC-AE at high resolutions, we re-use the tiling code by [undefr] to save training resources.
@@ 5.1 Image-to-Video Training
As illustrated in Figure 9, our approach first encodes image or video conditions using an autoencoder, then concatenates the encoded features with the original video latent representation. An extra channel is introduced to indicate the task type, increasing the number of channels from $k$ to $2k+1$ .
@@ 5.2 Image-to-Video Inference
We use classifier-free guidance for inference [undefl]. As our model is conditioned on both image and text, a straightforward approach is to use a single guidance scale:
@@
A further issue arises when using high image guidance, as it sometimes introduces flickering in generated videos. To mitigate this, we introduce a guidance oscillation technique [undefm] for image guidance. Taking 50-step sampling as an example, after the first 10 steps, we alternate the image guidance scale: during odd-numbered steps, we apply the original $g_{\textit{img}}$, while for even-numbered steps, we reduce it to 1. This helps balance stability and motion consistency.

### v2

@@ 1 Introduction
The past year has witnessed an explosion of video generation models. Since the emergence of OpenAI’s Sora [videoworldsimulators2024] in February 2024, a series of video generation models—either open-source [zheng2024open, lin2024open, ma2025step, kong2024hunyuanvideo, hong2022cogvideo] or proprietary [polyak2024moviegen, runwayml_gen3, luma_dream_machine]—have appeared at an unprecedented pace, each striving to achieve "Sora-level" generation quality. While the quality of generated videos continues to improve, there is a clear trend toward rapid growth in model size, data quantity, and computing resources. Following the success of scaling large language models (LLMs) [hoffmann2022training, kaplan2020scaling], researchers are now applying similar scaling principles [kong2024hunyuanvideo] to video generation, converging on similar model architectures and training techniques.
In this report, however, we show that a top-performing video generation model can be trained at a highly controlled cost, offering new insights into cost-effective training and efficiency optimization.. We build Open-Sora 2.0, a commercial-level video generation model trained for only \$200k. As shown in Table. 7, the training cost is 5-10 times lower than comparable models like MovieGen [polyak2024moviegen] and Step-Video-T2V [ma2025step] based on a fair comparison. This remarkable cost efficiency stems from our joint optimization of data curation, training strategy, and AI infrastructure—all of which are detailed in the following sections.
We show the human preference evaluation of Open-Sora 2.0 and other global leading video generation models in Figure 1. The models for comparison include proprietary models like Runway Gen-3 Alpha [runwayml_gen3], Luma Ray2 [luma_dream_machine] and open-source models like HunyuanVideo [kong2024hunyuanvideo] and Step-Video-T2V [ma2025step]. We evaluate these models in three important aspects: 1) visual quality, 2) prompt adherence, and 3) motion quality. Despite the low cost of our model, it outperforms these top-performing models in at least two aspects out of the three, demonstrating its strong potential for commercial deployment.
@@ 2.1.1 Preprocessing
Then, we detect continuous shots within the raw video and segment it into short clips based on the detection result. For shot detection, we apply the libavfilter from FFmpeg [FFmpeg] to calculate the scene scores, which quantify visual differences among frames. The shot boundaries are identified at points where the change in scene scores exceeds an empirically determined threshold.
@@ 2.1.2 Score Filtering
We assess the aesthetic quality of a sample using the CLIP+MLP aesthetic score predictor [schuhmann2021aesthetic]. For a video sample, we extract the first, middle, and last frame, compute their individual scores, and take the average as the final aesthetic score.
@@
The motion intensity of a video is measured using the VMAF motion score from FFmpeg’s libavfilter library [FFmpeg]. Videos with extremely low or excessively high motion scores are filtered out to maintain a balanced motion range.
@@
To evaluate image clarity, we apply the Laplacian operator from OpenCV [opencv_library]. A sample is considered blurry if the variance of the Laplacian image falls below a configurable threshold. For video samples, we extract five uniformly spaced frames and adopt a majority voting approach to determine blurriness.
@@
We detect text bounding boxes using PaddleOCR [paddleocr2020] and compute the total area of bounding boxes with a confidence score above 0.7. Images or videos containing excessive text are discarded to avoid unwanted overlays.
@@
We detect camera jitter using Shot Boundary Detection from the PySceneDetect library [pyscenedetect]. A video is identified as having camera jitter if the average frame-to-frame change exceeds a predefined threshold.
@@ 2.2 Data Annotation
For captioning, we utilize the open-source vision-language model LLaVA-Video [zhang2024video] to annotate 256px videos. We prompt the model to focus on six aspects for a detailed and comprehensive caption, which are 1) main subjects; 2) subjects’ actions; 3) background and environment; 4) lighting condition and atmosphere; 5) camera movement; 6) video style, such as realistic, cinematic, 3D, animation, and so on. For high-resolution 768px training data, we leverage a stronger proprietary model Qwen 2.5 Max [qwen25] to generate more accurate and semantically aligned captions. We find that Qwen 2.5 Max produces fewer hallucinations and delivers better semantic consistency than LLaVA-Video.
@@ 3 Model Architecture
Open Sora 1.2 [zheng2024open] $4\times 8\times 8$ 192 4 $\checkmark$ 0.161 27.504 0.756
StepVideo VAE [ma2025step] $8\times 16\times 16$ 96 64 $\checkmark$ 0.082 28.719 0.818
HunyuanVideo VAE [kong2024hunyuanvideo] $4\times 8\times 8$ 48 16 $\checkmark$ 0.046 30.240 0.856
@@
We begin by leveraging the open-source HunyuanVideo VAE [kong2024hunyuanvideo] as the initial autoencoder for our model. To further reduce both training and inference costs, we develop a video autoencoder with deep compression (Video DC-AE, named after [chen2024deep]) that improves efficiency while maintaining high reconstruction fidelity.
@@
[…]
chen2024deep introduce DC-AE as an effective approach to achieve higher downsampling ratios while maintaining reconstruction quality. While the original DC-AE is primarily optimized for image encoding, we extend its architecture to support video encoding by incorporating temporal compression mechanisms.
As shown in Figure 5(a), our Video DC-AE encoder consists of three residual blocks followed by three EfficientViT blocks [cai2023efficientvit], with the decoder adopting a symmetrical structure. The first five encoder blocks and the last five decoder blocks function as downsampling and upsampling layers, respectively.
@@
[…]
@@
We train Video DC-AE from scratch and evaluate its reconstruction quality, with results summarized in Table 1. Our model is compared against OpenSora 1.2 VAE, the recently open-sourced StepVideo VAE [ma2025step], and HunyuanVideo VAE [kong2024hunyuanvideo]. The results demonstrate that Video DC-AE achieves competitive performance, with only minor degradation in LPIPS [zhang2018unreasonable] scores compared to the best-performing HunyuanVideo VAE, while maintaining strong PSNR and SSIM scores.
@@ 3.2 DiT Architecture
To achieve higher video quality, we employ full attention to capture long-range dependencies effectively. After encoding the input through the autoencoder, we patchify the latent representations to enhance computational efficiency and improve model learning. Sana [xie2024sana] suggests that a patch size of 1 yields better training stability and finer details in video generation. With Video DC-AE’s high compression ratio, we can afford to reduce the patch size to 1 (i.e., no patch at all), whereas a patch size of 2 is still required when the HunyuanVideo autoencoder is used in our generation model.
Inspired by FLUX’s MMDiT [flux2024], we employ a hybrid transformer architecture that incorporates both dual-stream and single-stream processing blocks. In the dual-stream blocks, text and video information are processed separately to facilitate more effective feature extraction within each modality. Subsequently, single-stream blocks integrate these features to facilitate effective cross-modal interactions. To further enhance the model’s ability to capture spatial and temporal information, we apply 3D RoPE (Rotary Position Embedding) [su2024roformer], which extends traditional positional encoding to three-dimensional space, allowing the model to better represent motion dynamics across time.
For text encoding, we leverage T5-XXL [chung2024scaling] and CLIP-Large [Radford2021LearningTV], two high-capacity pretrained models known for their strong semantic understanding. T5-XXL captures complex textual semantics, while CLIP-Large improves alignment between text and visual concepts, leading to more accurate prompt adherence in video generation.
@@ 4.1 Efficient Training Strategy
Prior research [ma2024latte, hong2022cogvideo, zheng2024open] has demonstrated that pretraining on image datasets can significantly accelerate video model training. To avoid the high cost of training an image model from scratch, we leverage an open-source solution. Specifically, we adopt Flux [flux2024], a state-of-the-art text-to-image model with 11 billion parameters, which provides sufficient capacity to generate high-quality videos. We initialize our text-to-video model using Flux and empirically find that, despite it being a distilled model, this initialization is effective for further training.
@@
Inspired by the efficient image training strategies from PixArt [chen2023pixart], we hypothesize that high-quality video data can substantially enhance training efficiency. Consequently, we curate a high-quality subset from a large-scale dataset for low-resolution training. For high-resolution fine-tuning, we impose stricter selection criteria to ensure superior video quality.
@@
With our proposed approach, we successfully constrain the one-time training cost to \$200K, as detailed in Table 3. We estimate the training costs of Movie Gen [polyak2024moviegen] and Step-Video-T2V [ma2025step] based on publicly available information, as summarized in Figure 7. Our model achieves 5–10× lower training costs. We hope that these insights will contribute to reducing the training cost of high-quality video generation models in future research.
Figure 7: Training Cost Comparison of Different Video Generation Models. Values marked with * are estimated based on publicly available information. Our Open-Sora 2.0 is trained at 5–10× lower training cost.
@@ 4.2 Training Setting
Our training setup is largely based on our previous iteration, Open-Sora 1.2 [zheng2024open]. We adopt flow matching [lipman2022flow] as our primary training objective and utilize the AdamW [loshchilov2017decoupled] optimizer with $\beta$ values of (0.9, 0.999) and an $\epsilon$ value of $1\times 10^{-15}$. No weight decay is applied. The learning rate is set to $5\times 10^{-5}$ for the first 40k steps of Stages 1 and 2, followed by a decay to $3\times 10^{-5}$ for the final 45k steps. For Stage 3, it is reduced to $1\times 10^{-5}$. Gradient norm clipping is applied with a threshold of 1.
@@
Our flow matching approach is similar to that used in Stable Diffusion 3 [esser2024scaling]. We denote the video latent as $X_{0}$, and Gaussian noise $X_{1}\sim\mathcal{N}(0,1)$. The model $f_{\theta}$ takes as input an interpolated latent $X_{t}=(1-t)X_{0}+tX_{1}$ and is trained to predict the velocity component $X_{0}-X_{1}$. The corresponding loss function is formulated as:
@@
Following the methodology in Open-Sora 1.2 [zheng2024open], we adopt multi-bucket training to efficiently handle videos of varying frame counts, resolutions, and aspect ratios within the same batch. This strategy optimizes GPU utilization by dynamically assigning batch sizes based on video characteristics. The specific batch sizes used for different training configurations are detailed in Table 4 and Table 5.
@@ 4.3 High Compression Autoencoder Adaptation
For instance, in a 5-second, 24fps, 768px video, the HunyuanVideo VAE with a $4\times 8\times 8$ compression in combination of generation model’s patch size $1\times 2\times 2$ ($D_{\text{token}}=1024$) reduces the token count to 76K, whereas the Video DC-AE with $4\times 32\times 32$ compression and patch size of $1\times 1\times 1$ has a $D_{\text{token}}$ of 4096, effectively reduces the token count to 19K. As shown in Figure 8, this compression yields a 5.2×5.2$\times$5.2 × speedup in training throughput. Additionally, inference speeds improve by over 10×10$\times$10 ×, making it a highly efficient approach.
@@
With the growing interest in video generation models, we observe that despite variations in architecture, downsampling ratios, number of channels, and training losses, video autoencoders trained with similar information downsampling ratios exhibit comparable performance (evaluated at 256px resolution) [chen2024deep]. We define the information downsampling ratio, $D_{\text{info}}$ as follows:
@@
However, training a deep video compression network is challenging. Prior work [chen2024deep] identifies gradient propagation issues in downsample and upsample blocks. To address this, they introduce residual connections in these blocks using pixel shuffle and unshuffle operations. We verify that training the HunyuanVideo VAE architecture with residual connections improves performance. Building on this, we incorporate temporal downsampling into DC-AE to construct Video DC-AE.
@@
[…]
@@
    Latent Space Distillation: After training Video DC-AE, we apply a distillation loss to align the third-layer latents with DINOv2 [oquab2023dinov2], leveraging their similar latent shapes.
@@
[…]
@@ 5.1 Image-to-Video Training
As illustrated in Figure 10, our approach first encodes image or video conditions using an autoencoder, then concatenates the encoded features with the original video latent representation. An extra channel is introduced to indicate the task type, increasing the number of channels from $k$ to $12k+$ 1 .
@@ 5.2 Image-to-Video Inference
We use classifier-free guidance for inference [ho2022classifier]. As our model is conditioned on both image and text, a straightforward approach is to use a single guidance scale:
@@
A further issue arises when using high image guidance, as it sometimes introduces flickering in generated videos. To mitigate this, we introduce a guidance oscillation technique [ho2022imagen] for image guidance. Taking 50-step sampling as an example, after the first 10 steps, we alternate the image guidance scale: during odd-numbered steps, we apply the original $g_{\textit{img}}$, while for even-numbered steps, we reduce it to 1. This helps balance stability and motion consistency.